OpenStudy (anonymous):

what is the solution l 2x-1 l=1

OpenStudy (austinl):

\(|2x-1|=1\) This is the same as, \(2x-1=1\) \(-(2x-1)=1\) Solve each of those equations for x, and that should give you your two solutions!